PLease help to graph on a number line
x^2 - 5x > or equal to -6; D = {Reals}

Answer by stanbon(48516)

x^2 - 5x > or equal to -6
x^2-5x+6 >=0
(x-6)(x+1) >=0
Draw a number line.
Mark 6 and -1 on the line
Use a test value from each interval to see where the solution set is.
If x=-10, (-16)(-9)>=0 True; so the interval (-inf,-1) is part of the solution
If x=0, (-6)(1)>=0 is False; so no solutions in interval (-1,6)
If x=10, (4)(11)>0 is True; so the interval (6,inf) is part of the solution.
FINAL ANSWER: (-inf,-1) or (6,inf)
Cheers,
Stan H.